OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (-3, -1) with a slope of 4.

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1\\ &(-3\quad ,&-1) \end{array} \\\quad \\ slope = m= 4 \\ \quad \\ y-y_1=m(x-x_1)\quad \textit{plug in your values and solve for "y"}\) to get the equation